{-3x + y = 4
{-x + y = -2

Answer by pannalal(9)

Given Equations:
[1] -3x + y = 4
[2] -x + y = -2
-3x + y = 4 ..........(1)
-x + y = -2 ..........(2)
On Subtracting Eq(2) from Eq(1), we get:
-2x = 6 ..............(3)
From Eq(3), we get:
x = -3 ...............(4)
On Substituting value of x in Eq(1), we get:
-3*(-3) + y = 4
y = -5 ...............(5)
Summary:
x = -3
y = -5